40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Tue Jun 22 14:35:40 2021, aaron, shruti, DailyProgress, Control System, slow temperature control 7x
    Reply  Wed Jun 23 17:53:21 2021, aaron, DailyProgress, Control System, slow temperature control 
       Reply  Thu Jun 24 12:26:56 2021, shruti, DailyProgress, Control System, slow temperature control 
          Reply  Thu Jun 24 14:37:12 2021, aaron, DailyProgress, Control System, slow temperature control, more transfer functions 
             Reply  Mon Jun 28 10:12:28 2021, shruti, DailyProgress, Control System, glitches, transfer function glitches.pngOLTF.pdfSetup.pdfData.zip
                Reply  Mon Jun 28 13:16:07 2021, aaron, DailyProgress, Control System, glitches 
                   Reply  Tue Jun 29 08:53:02 2021, Chris, HowTo, Control System, acquiring slow channels 
Message ID: 2761     Entry time: Tue Jun 22 14:35:40 2021     Reply to this: 2763
Author: aaron, shruti 
Type: DailyProgress 
Category: Control System 
Subject: slow temperature control 

[aaron, shruti]

I observed an unexpected behavior this afternoon that I still can't explain. I managed to get the cavity locked using the LB box servo in LFGL mode. When I turned off the servo box by switching to 'lock off' mode, the cavity maintained lock and the PDH error signal was passed through to the current modulation point. Only the LB servo was driving the modulation point, and the temperature tuning was also off. My understanding from the LB manual is that in 'lock off' mode, no control signal is summed into the output signal... so why was the PDH error signal passed through?

Later, we started controlling the temperature of the laser diode using some slow epics channels.

  1. Lock the cavity using the LB box (in LFGL mode) to modulate the current
  2. Turn on the python PID loop. The script and configuration we used are in controls@spirou:~/cryo_lab/scripts/temp_control/.
    • Operated the temperature control loop at 5 Hz in pure integrator mode, with K_I = 0.1-0.5       
    • I have the PDH setpoint at 0, and checked that any DC offset on the ADC is small (< 10 counts)
  3. Next, we tuned the input and sweep offsets of the LB box to optimize our dynamic range. We need the temperature control loop on for this operation, to avoid railing the current modulation input at the laser driver
    • first, tuned the LB box input offset until the error monitor was centered on 0 V
    • Next, tuned the LB box sweep offset until the current control was centered on 0 V
  4. With the low frequency gain limited to 20 dB, the PDH error signal was wandering at a few Hz. We turned up the gain limit to 40 dB, though we could have increased the gain of the temperature PID until it took over at those frequencies.

We observed a 100 kHz oscillation in the noise spectrum after this procedure. We weren't able to change the oscillation by tuning the laser current (within 10 mA) or servo gain (while maintaining lock).

We measured the open loop transfer function of only the LB servo box (feeding back its output to -B), and didn't see a feature at 100 kHz or an oscillation in the noise spectrum. We measured the transfer function in both 'lock on' and 'LFGL' modes. We did observe a broad peak near 4.5 MHz in the noise measured at the LB error monitor (attachment 1, 2. The sharp peaks are artifacts from the Moku present even with no input connected).

Data are available in the ligo.wbridge google drive. Attachment 7 shows the broad 100 kHz oscillation on the PDH error signal in purple.

 

Attachment 3: Updated OLTFs with LB1005 measured separately

We measured the OLTF of the LB1005 servo by feeding back to itself with the setting 'LFGL' (Low Frequency Gain Limit) set to 40 dB, INT (pure integrator until it hits LFGL), and gain of 5.1 at 300 kHz.

 I (Shruti) think the gain setting was slightly different today which makes the green curve slightly higher in magnitude than the orange curve, but otherwise it seems to track it and do nothing strange. The orange curve is the TF of the LB1005 derived when the servo is used to lock the laser to the PSOMA cavity as in elog 2759. The phase for both measurements also seems to be the same up to 1 MHz.

Data for LB1005 in Attachment 4 and remaining data used for the plots in Attachments 3 and 4 of elog 2759.

 

Attachment 5: Measured OLTFs again with new settings

After these changes we measured the loop TFs again. It is strange that the phase of the open loop first increases and then decreases.

Also there is a weird dip at 80 kHz (right above the UGF) in both the Plant TF and the full open loop TF.

The LB1005 measured separately and in the full closed loop differ by ~4dB, the full loop setting resulting in the lower curve, even with the same gain setting. Otherwise the two run almost parallel, at least below 1 MHz.

The data for this is in Attachment 6.

Attachment 1: MokuSpectrumAnalyzerData_20210622_152905_1LB_2ctrl_LB1005only_Screenshot.png  1.688 MB  Uploaded Tue Jun 22 16:38:54 2021  | Hide | Hide all
MokuSpectrumAnalyzerData_20210622_152905_1LB_2ctrl_LB1005only_Screenshot.png
Attachment 2: MokuSpectrumAnalyzerData_20210622_153041_1LB_2ctrl_LB1005only_Screenshot.png  1.737 MB  Uploaded Tue Jun 22 16:39:13 2021  | Hide | Hide all
MokuSpectrumAnalyzerData_20210622_153041_1LB_2ctrl_LB1005only_Screenshot.png
Attachment 3: OLTF_wLBonly.pdf  47 kB  Uploaded Tue Jun 22 16:44:58 2021  | Hide | Hide all
OLTF_wLBonly.pdf
Attachment 4: LB1005TF.zip  1.593 MB  Uploaded Tue Jun 22 16:50:10 2021
Attachment 5: OLTF_wLB_20210622.pdf  43 kB  Uploaded Tue Jun 22 17:40:35 2021  | Hide | Hide all
OLTF_wLB_20210622.pdf
Attachment 6: 20210622LoopTFs.zip  2.712 MB  Uploaded Tue Jun 22 17:49:44 2021
Attachment 7: 65DBE9F2-84F8-4BA1-9A8B-4CF90D1ADE8A.jpeg  2.953 MB  Uploaded Wed Jun 23 10:57:27 2021  | Hide | Hide all
65DBE9F2-84F8-4BA1-9A8B-4CF90D1ADE8A.jpeg
ELOG V3.1.3-